Put the quadratic into vertex form and state the coordinates of the vertex.
y=x^2-14x

Answer by ikleyn(52777)

.

Complete the square

    y = x^2 - 14x = (x^2 - 14x + 49) - 49 = (x-7)^2 - 49.


So, the vertex form is 

    y = (x-7)^2 - 49.


Coordinates of the vertex are x= 7, y= -49.     ANSWER

Solved.

--------------

On completing the square see the lesson
    - HOW TO complete the square - Learning by examples
in this site.

On finding the vertex of a parabola, see the lessons
    - HOW TO complete the square to find the vertex of a parabola
    - Briefly on finding the vertex of a parabola
in this site.